### v1

# The ECME Algorithm Using Factor Analysis for DOA Estimation in Nonuniform Noise

Abstract
Maximum likelihood factor analysis has been used for direction of arrival estimation in unknown nonuniform noise and some iterative approaches have been developed. In particular, the Factor Analysis for Anisotropic Noise (FAAN) method proposed by Stoica and Babu has excellent convergence properties. In this letter, the Expectation/Conditional Maximization Either (ECME) algorithm, an extension of the expectation-maximization algorithm, is designed, which has almost the same computational complexity at each iteration as the FAAN method. However, numerical results show that the ECME algorithm yields faster stable convergence and is computationally more efficient.
Index Terms:
Array signal processing, direction of arrival (DOA) estimation, expectation-maximization (EM) algorithm, factor analysis, maximum likelihood.
I Introduction
Factor analysis (FA) has been a tool in psychology for over one hundred years [1] and its purpose is to extract or estimate the low-rank covariance matrix from a measurement covariance matrix. To this end, several techniques have been developed, but the only one available, which provides efficient and consistent estimators, is maximum likelihood FA (MLFA) [2]. Unfortunately, these estimators cannot be expressed in closed form, which leads to a variety of iterative approaches [3]. In particular, one of the most promising is the Expectation-Maximization (EM) algorithm [4].
The EM algorithm regards observations or measurements as the incomplete data and designs the complete data containing missing (latent or unobserved) data. The complete data is so manipulable that its log-likelihood function (LLF) is more tractable than the incomplete-data or actual LLF. Hence, the EM algorithm uses the complete-data LLF to efficiently update the estimates of parameters, which can still increase the actual likelihood. To be specific, each iteration of the EM algorithm consists of an expectation step (E-step) and a maximization step (M-step). However, the M-step requires maximizing the complete-data LLF, which causes two shortcomings: 1) slow convergence, and 2) difficult maximization in numerous complex cases. To make up for both shortcomings, the Expectation/Conditional Maximization Either (ECME) algorithm, an extension of the EM algorithm, has been proposed [5] and then applied to MLFA [6]. However, a conditional M-step (CM-step) at each iteration utilizes Newton-Raphson to update the estimates of partial parameters by the actual LLF and thus is complex as well as unstable.
In addition to psychometrics, econometrics, statistics and so on, FA recently finds applications in array signal processing, especially including direction of arrival (DOA) estimation in unknown nonuniform noise [7, 8]. The nonuniform noise assumption can let the signal received at an array be modeled by FA so that MLFA can be utilized to estimate the signal subspace for the subspace based DOA estimation [9]. For this purpose, an iterative approach, called Iterative ML Subspace Estimation (IMLSE), is proposed in [7]. This method sequentially updates the estimates of parameters at each iteration by explicit formulas and is computationally attractive. However, Stoica and Babu have shown that the IMLSE method cannot guarantee its monotonicity and may fail to converge [8]. Then, they propose a new iterative approach, called Factor Analysis for Anisotropic Noise (FAAN), with excellent convergence properties.
In this letter, the ECME algorithm is designed again for MLFA, which has almost the same computational complexity at each iteration as the FAAN method. However, numerical results show that the ECME algorithm yields faster stable convergence and is computationally more efficient. Importantly, the signal subspace estimated by the ECME algorithm can be used for the subspace based DOA estimation in unknown nonuniform noise.
Notations: $(\cdot)^{T}$ and $(\cdot)^{H}$ mean the transpose and conjugate transpose operators, respectively. $\mathrm{Det}(\boldsymbol{\Sigma})$, $\mathrm{Tr}\{\boldsymbol{\Sigma}\}$, $\boldsymbol{\Sigma}^{1/2}$, and $\boldsymbol{\Sigma}^{-1}$ stands for the determinant, trace, square root, and inverse of square matrix $\boldsymbol{\Sigma}$. $\Sigma_{i,j}$ is the element in the $i$th row and $j$th column of $\boldsymbol{\Sigma}$. $\boldsymbol{\Sigma}\succ\mathbf{0}_{N}$ and $\boldsymbol{\Sigma}\succeq\mathbf{0}_{N}$ denote that the $N\times N$ matrix $\boldsymbol{\Sigma}$ is positive definite and positive semi-definite, respectively. $\mathbf{I}_{N}$ is the identity matrix of order $N$ and $\jmath$ is the imaginary unit. $\mathcal{C}(\boldsymbol{\Sigma})$ denotes the column space of $\boldsymbol{\Sigma}$. $(\cdot)^{(k)}$ is an iterative value obtained at the $k$th iteration.
II Problem Description
We consider a uniform linear array (ULA) consisting of $N$ isotropic sensors for analytical simplicity. $M<N$ narrow-band source signals are assumed to impinge upon the array from far field and the distance between arbitrary two adjacent sensors is the half-wavelength of the source signals. Thus, the composite signal received at this ULA can be expressed as
$\mathbf{y}(t)=\sum_{m=1}^{M}\mathbf{a}(\theta_{m})s_{m}(t)+\mathbf{v}(t)=\mathbf{A}\mathbf{s}(t)+\mathbf{v}(t),$ (1)
where $\theta_{m}\in(0,\pi)$ represents the DOA of the $m$th source, $\mathbf{a}(\theta_{m})=[1~\cdots~e^{-(N-1)\jmath\pi\cos(\theta_{m})}]^{T}\in\mathbb{C}^{N}$ is the steering vector of the $m$th source, and $\mathbf{A}=[\mathbf{a}(\theta_{1})~\cdots~\mathbf{a}(\theta_{M})]$ is the array manifold matrix. Additionally, $\mathbf{s}(t)=[s_{1}(t)~\cdots~s_{M}(t)]^{T}\in\mathbb{C}^{M}$ is the source signal vector, $\mathbf{v}(t)$ means a spatially nonuniform Gaussian white noise vector with mean zero and covariance $\mathbf{Q}=\mathrm{diag}\{\sigma^{2}_{1},\dots,\sigma^{2}_{N}\}\succ\mathbf{0}_{N}$, i.e., $\mathbf{v}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{Q})$.
When the stochastic source signal model is adopted, $\mathbf{s}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{P})$, where $\mathbf{P}\succeq\mathbf{0}_{M}$ is the source covariance matrix. In (1), $\mathbf{s}(t)$ and $\mathbf{v}(t)$ are assumed to be mutually uncorrelated, which leads to $\mathbf{y}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{C})$ with $\mathbf{C}=\mathbf{A}\mathbf{P}\mathbf{A}^{H}+\mathbf{Q}$ [9]. In practice, we first sample $\mathbf{y}(t)$ and obtain independent and identically distributed snapshots as follows:
$\mathbf{y}(t)=\mathbf{A}\mathbf{s}(t)+\mathbf{v}(t),t=1,\dots,L$ (2)
where $L$ is the number of snapshots. Then, we obtain the DOA estimates $\hat{\theta}_{m}$’s by processing the snapshots $\mathbf{y}(t)$’s.
From the perspective of ML direction finding [10], we need to write the LLF of $\mathbf{Y}$ with $\mathbf{Y}=[\mathbf{y}(1)~\cdots~\mathbf{y}(L)]$ as
$l(\boldsymbol{\Phi})$ $=$ $\ln p(\mathbf{Y};\boldsymbol{\Phi})={\sum}_{t=1}^{L}\ln p\big{(}\mathbf{y}(t);\boldsymbol{\Phi}\big{)}$
$=$ $-LN\ln(\pi)-L\Big{(}\ln\mathrm{Det}(\mathbf{C})+\mathrm{Tr}\{\hat{\mathbf{R}}\mathbf{C}^{-1}\}\Big{)}$
where $\boldsymbol{\Phi}=(\boldsymbol{\theta},\mathbf{P},\mathbf{Q})$ with $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{M})$, $p(\cdot)$ is the probability density function, and $\hat{\mathbf{R}}=(1/L)\sum_{t=1}^{L}\mathbf{y}(t)\mathbf{y}^{H}(t)\succ\mathbf{0}_{N}$ [^1] In this letter, we assume $\hat{\mathbf{R}}\succ\mathbf{0}_{N}$, which requires $L\geq N$.. Importantly, we need to solve the equivalent ML direction finding problem [11, 12]:
$\min_{\boldsymbol{\theta}\in(0,\pi)^{M},\mathbf{P}\succeq\mathbf{0}_{M},\mathbf{Q}\succ\mathbf{0}_{N}}\ln\mathrm{Det}(\mathbf{C})+\mathrm{Tr}\{\hat{\mathbf{R}}\mathbf{C}^{-1}\}.$ (3)
In order to solve problem (3), the ECME algorithm has been designed in the previous work [13]. In the next section, we will design the ECME algorithm using FA for the subspace based DOA estimation.
III MLFA Based Subspace Estimation
The ML direction finding problem (3) is intractable since $\mathbf{C}=\mathbf{APA}^{H}+\mathbf{Q}$ has excessive constraints on free parameters (unknowns) to be estimated. To alleviate this issue, we should release as many constraints as possible and introduce as many free parameters as possible. Thus, we aim not to solve problem (3) in this letter.
III-A MLFA Estimation Problem
It is well-known that the subspace based DOA estimation only requires the signal subspace $\mathcal{C}(\mathbf{A})$ [9]. As a result, let $\mathbf{C}$ be the population covariance matrix in FA [3]
$\mathbf{C}=\mathbf{S}\mathbf{S}^{H}+\mathbf{Q},$ (4)
where all elements in $\mathbf{S}\in\mathbb{C}^{N\times M}$ are free parameters (factor loadings) and hence $\mathbf{SS}^{H}$ must be able to equal any $\mathbf{APA}^{H}$. Here, we also use $\mathbf{S}$ to estimate $\mathcal{C}(\mathbf{A})$. Eq. (4) makes problem (3) become the classic MLFA estimation problem [2]:
$\min_{\mathbf{S},\mathbf{Q}\succ\mathbf{0}_{N}}\ln\mathrm{Det}(\mathbf{C})+\mathrm{Tr}\{\hat{\mathbf{R}}\mathbf{C}^{-1}\}.$ (5)
For solving problem (5), the IMLSE method utilizes a fixed-point iteration and sequentially (not simultaneously) updates the estimates of $\mathbf{S}$ and $\mathbf{Q}$ at each iteration by explicit formulas [7]. However, Stoica and Babu have shown that the IMLSE method cannot guarantee its monotonicity and may fail to converge [8]. Hence, they propose the FAAN method with excellent convergence properties. We outline the FAAN method in the next subsection.
III-B FAAN Method
When a direct minimization over all free parameters is very difficult, an iterative (coordinate descent) approach or alternating minimization can always be adopted [14]. As mentioned before, the FAAN method, composed of two steps at each iteration, has been proposed for problem (5). Specifically, since $\mathbf{C}=\mathbf{SS}^{H}+\mathbf{Q}$ makes updating the estimates of $\mathbf{S}$ and $\mathbf{Q}$ difficult, a new parametrization of $\mathbf{C}$ is designed as
$\mathbf{C}=\mathbf{Q}^{1/2}(\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{H}+\mathbf{I}_{N})\mathbf{Q}^{1/2},$ (6)
where $\boldsymbol{\Lambda}\succeq\mathbf{0}_{M}$ and $\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{H}$ is the eigen-decomposition of $\mathbf{Q}^{-1/2}\mathbf{SS}^{H}\mathbf{Q}^{-1/2}$. Thus, problem (5) is changed to
$\min_{\mathbf{U},\boldsymbol{\Lambda}\succeq\mathbf{0}_{M},\mathbf{Q}\succ\mathbf{0}_{N}}\ln\mathrm{Det}(\mathbf{Q})+\ln\mathrm{Det}(\boldsymbol{\Lambda}+\mathbf{I}_{M})+$
$\mathrm{Tr}\big{\{}\mathbf{Q}^{-1/2}\hat{\mathbf{R}}\mathbf{Q}^{-1/2}(\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{H}+\mathbf{I}_{N})^{-1}\big{\}}.$ (7)
According to (III-B), the first step at the $k$th iteration of the FAAN method simultaneously obtains $\mathbf{U}^{(k)}$ and $\boldsymbol{\Lambda}^{(k)}$ by
$\min_{\mathbf{U},\boldsymbol{\Lambda}\succeq\mathbf{0}_{M}}\ln\mathrm{Det}(\boldsymbol{\Lambda}+\mathbf{I}_{M})+\mathrm{Tr}\big{\{}\tilde{\mathbf{R}}^{(k)}(\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{H}+\mathbf{I}_{N})^{-1}\big{\}},$ (8)
where $\tilde{\mathbf{R}}^{(k)}=[\mathbf{Q}^{(k-1)}]^{-1/2}\hat{\mathbf{R}}[\mathbf{Q}^{(k-1)}]^{-1/2}$. This leads to
$\mathbf{U}^{(k)}$ $=$ $\mathbf{B}^{(k)}\in\mathbb{C}^{N\times M},$ (9a)
$\Lambda_{m,m}^{(k)}$ $=$ $(\lambda_{m}^{(k)}-1)_{+},m=1,\dots,M$ (9b)
where the $\lambda_{m}^{(k)}$’s are the eigenvalues of $\tilde{\mathbf{R}}^{(k)}$ and $\lambda_{1}^{(k)}\geq\cdots\geq\lambda_{N}^{(k)}$, $\mathbf{B}^{(k)}$ is composed of the $M$ principal eigenvectors. Moreover, the operator $(\cdot)_{+}$ replaces the negative value with zero and leaves the nonnegative value unchanged.
The second step at the $k$th iteration of the FAAN method sequentially (not simultaneously) obtains $\sigma_{1}^{(k)},\dots,\sigma_{N}^{(k)}$ by
$\min_{\sigma_{n}>0}\ln\mathrm{Det}(\mathbf{Q})+\mathrm{Tr}\{\mathbf{Q}^{-1/2}\hat{{\mathbf{R}}}\mathbf{Q}^{-1/2}\boldsymbol{\Gamma}^{(k)}\},$ (10)
where $\boldsymbol{\Gamma}^{(k)}=\big{(}\mathbf{U}^{(k)}\boldsymbol{\Lambda}^{(k)}[\mathbf{U}^{(k)}]^{H}+\mathbf{I}_{N}\big{)}^{-1}$ [^2] We do not adopt $\boldsymbol{\Gamma}^{(k)}=\mathbf{I}_{N}+\mathbf{U}^{(k)}\big{(}[\boldsymbol{\Lambda}^{(k)}]^{-1}-\mathbf{I}_{M}\big{)}[\mathbf{U}^{(k)}]^{H}$ since numerical results show its low accuracy.. This leads to
$\sigma_{n}=\frac{1}{2}\Big{(}b_{n}^{(k)}+\sqrt{[b_{n}^{(k)}]^{2}+4c_{n}^{(k)}}\Big{)}>0,n=1,\dots,N$ (11)
where $\Re(\cdot)$ denotes the real part of the complex value, $b_{n}^{(k)}=\sum_{i\neq n}\Re\{\hat{R}_{i,n}\Gamma_{i,n}^{(k)}\}/\sigma_{i}$ and $c_{n}^{(k)}=\hat{R}_{n,n}\Gamma_{n,n}^{(k)}>0$. The details of the FAAN method are in Algorithm 1.
Algorithm 1 FAAN Method
1: Initialize $\mathbf{Q}^{(0)}$, $K_{1}$, $K_{2}$, and $k=1$.
2:while $k\leq K_{1}$ do
3:  Obtain $\mathbf{U}^{(k)}$ and $\boldsymbol{\Lambda}^{(k)}$ using (9).
4:  Obtain $\mathbf{Q}^{(k)}$ using (11) $K_{2}$ times.
5:  $k=k+1$.
6:end while
7: Let $\hat{\mathbf{S}}=[\mathbf{Q}^{(K_{1})}]^{1/2}\mathbf{U}^{(K_{1})}[\boldsymbol{\Lambda}^{(K_{1})}]^{1/2}$ and output $\hat{\mathbf{S}}$.
IV ECME Algorithm
The ECME algorithm has been designed for problem (5) in [6] and yields rapider convergence than the EM algorithm in [15]. However, a CM-step at each iteration utilizes Newton-Raphson to update the estimate of $\mathbf{Q}$ by the actual or negative LLF $f(\mathbf{S},\mathbf{Q})=\ln\mathrm{Det}(\mathbf{C})+\mathrm{Tr}\{\hat{\mathbf{R}}\mathbf{C}^{-1}\}$ in (5), which leads to that this step is complex and the monotonicity cannot be guaranteed. In this section, we again design the ECME algorithm with excellent convergence.
IV-A Algorithm Procedure
For problem (5), we do not adopt the complete data $(\mathbf{Y},\mathbf{X})$ with $\mathbf{X}=\big{[}\mathbf{x}(1)~\cdots~\mathbf{x}(L)\big{]}$ in [6, 15] and adopt the simpler complete data $(\mathbf{X},\mathbf{V})$ with $\mathbf{V}=\big{[}\mathbf{v}(1)~\cdots~\mathbf{v}(L)\big{]}$ [^3] Numerical results show that the ECME algorithm with $(\mathbf{X},\mathbf{V})$ yields the same convergence as that with $(\mathbf{Y},\mathbf{X})$. However, the ECME algorithm with $(\mathbf{X},\mathbf{V})$ needs fewer computations at each iteration.. Thus, we rewrite the snapshots or measurements $\mathbf{Y}$ in (2) by
$\mathbf{y}(t)=\mathbf{S}\mathbf{x}(t)+\mathbf{v}(t),t=1,\dots,L$ (12)
where $\mathbf{x}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{I}_{M})$ contains $M$ common factors and is uncorrelated with $\mathbf{v}(t)$. Then, the complete-data LLF is
$l(\mathbf{X},\mathbf{V};\mathbf{Q})$ $=$ $\ln p(\mathbf{X})+\ln p(\mathbf{V};\mathbf{Q})$ (13)
$=$ $c-L\Big{(}\ln\mathrm{Det}(\mathbf{Q})+\mathrm{Tr}\{\hat{\mathbf{R}}_{v}\mathbf{Q}^{-1}\}\Big{)}$
where $c$ is a term unrelated to any free parameters in $(\mathbf{S},\mathbf{Q})$, $\hat{\mathbf{R}}_{v}=(1/L)\sum_{t=1}^{L}\mathbf{v}(t)\mathbf{v}^{H}(t)$. With (13), we can construct the ECME algorithm, whose E-step and two CM-steps at the $k$th iteration are derived below.
IV-A1 First CM-step
To avoid providing $\mathbf{S}^{(0)}$, let this step only obtain $\mathbf{S}^{(k)}$ by minimizing $f(\mathbf{S},\mathbf{Q})$ with respect to $\mathbf{S}$, i.e.,
$\min_{\mathbf{S}}\ln\mathrm{Det}\big{(}\mathbf{SS}^{H}+\mathbf{Q}^{(k-1)}\big{)}+\mathrm{Tr}\big{\{}\hat{\mathbf{R}}(\mathbf{SS}^{H}+\mathbf{Q}^{(k-1)})^{-1}\big{\}}.$ (14)
This problem is very classic in MLFA and its classic solution is [2, 3]
$\mathbf{S}^{(k)}=[\mathbf{Q}^{(k-1)}]^{1/2}\mathbf{B}^{(k)}[\boldsymbol{\Lambda}^{(k)}]^{1/2}\in\mathbb{C}^{N\times M},$ (15)
where $\mathbf{B}^{(k)}$ and $\boldsymbol{\Lambda}^{(k)}$ are from (9).
IV-A2 E-step
Given $\mathbf{Y}$, $\mathbf{S}^{(k)}$, and $\mathbf{Q}^{(k-1)}$, this step finds the expectation of (13) by
$l(\mathbf{Q};\boldsymbol{\Omega}^{(k)})$ $=$ $\mathrm{E}\big{\{}l(\mathbf{X},\mathbf{V};\mathbf{Q})\mid\mathbf{Y};\boldsymbol{\Omega}^{(k)}\}$ (16)
$=$ $c-L\Big{(}\ln\mathrm{Det}(\mathbf{Q})+\mathrm{Tr}\{\hat{\mathbf{R}}_{v}^{(k)}\mathbf{Q}^{-1}\}\Big{)}$
where $\boldsymbol{\Omega}^{(k)}=(\mathbf{S}^{(k)},\mathbf{Q}^{(k-1)})$, ${}\mathrm{E}\{\cdot\}$ and ${}\mathrm{D}\{\cdot\}$ means expectation and covariance, respectively. Moreover,
$\hat{\mathbf{R}}^{(k)}_{v}=\mathrm{E}\big{\{}\hat{\mathbf{R}}_{v}\mid\mathbf{Y};\boldsymbol{\Omega}^{(k)}\big{\}}=[\boldsymbol{\delta}^{(k)}]^{H}\hat{\mathbf{R}}\boldsymbol{\delta}^{(k)}+\boldsymbol{\Delta}^{(k)}\succ\mathbf{0}_{N}$
where $\mathbf{C}^{(k)}=\mathbf{S}^{(k)}[\mathbf{S}^{(k)}]^{H}+\mathbf{Q}^{(k-1)}$, $\boldsymbol{\delta}^{(k)}=[\mathbf{C}^{(k)}]^{-1}\mathbf{Q}^{(k-1)}$, and $\boldsymbol{\Delta}^{(k)}=\mathbf{Q}^{(k-1)}-\mathbf{Q}^{(k-1)}\boldsymbol{\delta}^{(k)}\succeq\mathbf{0}_{N}$ [^4] We do not adopt $[\mathbf{C}^{(k)}]^{-1}=\mathbf{q}^{(k)}-\mathbf{q}^{(k)}\mathbf{S}^{(k)}(\boldsymbol{\Lambda}^{(k)}+\mathbf{I}_{M})^{-1}[\mathbf{S}^{(k)}]^{H}\mathbf{q}^{(k)}$ with $\mathbf{q}^{(k)}=[\mathbf{Q}^{(k-1)}]^{-1}$ since numerical results show its low accuracy.. The conditional distribution of $\mathbf{x}(t)$ can be derived from [16], and
$\mathrm{E}\big{\{}\mathbf{v}(t)\mid\mathbf{Y};\boldsymbol{\Omega}^{(k)}\big{\}}$ $=$ $[\boldsymbol{\delta}^{(k)}]^{H}\mathbf{y}(t),$
$\mathrm{D}\big{\{}\mathbf{v}(t)\mid\mathbf{Y};\boldsymbol{\Omega}^{(k)}\big{\}}$ $=$ $\boldsymbol{\Delta}^{(k)}\succeq\mathbf{0}_{N}.$
IV-A3 Second CM-step
This step only obtains $\mathbf{Q}^{(k)}$ by maximizing $l(\mathbf{Q};\boldsymbol{\Omega}^{(k)})$ in (16) with respect to $\mathbf{Q}$, i.e.,
$\min_{\mathbf{Q}\succ\mathbf{0}_{N}}\ln\mathrm{Det}(\mathbf{Q})+\mathrm{Tr}\big{\{}\hat{\mathbf{R}}_{v}^{(k)}\mathbf{Q}^{-1}\}.$ (17)
This problem can be decomposed into the $N$ parallel subproblems:
$\Big{(}\min_{\sigma_{n}^{2}>0}\ln(\sigma_{n}^{2})+[\hat{\mathbf{R}}_{v}^{(k)}]_{n,n}/\sigma^{2}_{n}\Big{)}~~n=1,\dots,N.$ (18)
The solutions are
$[\sigma^{2}_{n}]^{(k)}$ $=$ $[\hat{\mathbf{R}}_{v}^{(k)}]_{n,n}>0,n=1,\dots,N,$
$\Rightarrow\mathbf{Q}^{(k)}$ $=$ $\mathrm{diag}\big{(}\hat{\mathbf{R}}_{v}^{(k)}\big{)}\succ\mathbf{0}_{N},$ (19)
where $\mathrm{diag}(\cdot)$ is the matrix composed of the diagonal part. The details of the ECME algorithm are in Algorithm 2.
Algorithm 2 ECME Algorithm
1: Initialize $\mathbf{Q}^{(0)}$, $K$, and $k=1$.
2:while $k\leq K$ do
3:  Obtain $\mathbf{S}^{(k)}[\mathbf{S}^{(k)}]^{H}$ using (15).
4:  Obtain $\mathbf{Q}^{(k)}$ using (IV-A3).
5:  $k=k+1$.
6:end while
7: Let $\hat{\mathbf{S}}=\mathbf{S}^{(K)}$ and output $\hat{\mathbf{S}}$.
IV-B Convergence and Complexity
From [5, 17], we know that the ECME algorithm satisfies certain regularity conditions and its sequence $\{(\mathbf{S}^{(k)},\mathbf{Q}^{(k)})\}$ stably converges to a stationary point of $f(\mathbf{S},\mathbf{Q})$. Moreover, Eq. (15) and Eq. (IV-A3) guarantee that the solution provided by the ECME algorithm is feasible for problem (5).
By comparing Algorithm 1 and Algorithm 2, we can find that at the $k$th iteration, the computational complexity of Algorithm 1 lies in the eigen-decomposition of $\tilde{\mathbf{R}}^{(k)}\in\mathbb{C}^{N\times N}$ in (9) and the inverse of $\big{(}\mathbf{U}^{(k)}\boldsymbol{\Lambda}^{(k)}[\mathbf{U}^{(k)}]^{H}+\mathbf{I}_{N}\big{)}\in\mathbb{C}^{N\times N}$ in (11). Also, the computational complexity of Algorithm 2 lies in the eigen-decomposition of $\tilde{\mathbf{R}}^{(k)}\in\mathbb{C}^{N\times N}$ in (15) and the inverse of $\mathbf{C}^{(k)}\in\mathbb{C}^{N\times N}$ in (IV-A3). Thus, the ECME algorithm has almost the same computational complexity at each iteration as the FAAN method.
V Numerical Results
Simulation results are used to show and compare Algorithm 1 and Algorithm 2. Let $N=6$, $M=2$, $\mathbf{P}=10\times\mathbf{I}_{M}$, $\mathbf{Q}=\mathrm{diag}\{10,2,3,2,1,3\}$, $\theta_{1}=60^{\circ}$, $\theta_{2}=120^{\circ}$, and $K_{2}=100$. Moreover, $\mathbf{Q}^{(0)}=\mathbf{I}_{N}$.
Without loss of generality, Fig. 1 illustrates the convergence performances of both algorithms given the same snapshots. Here, $L=100$. We can clearly see that the ECME algorithm yields faster convergence than the FAAN method. Because the ECME algorithm has almost the same computational complexity at each iteration as the FAAN method, it is computationally more efficient.
Figure 1: Convergence comparison of the FAAN method and the ECME algorithm.
After obtaining $\hat{\mathbf{S}}$ or the signal subspace estimate $\mathcal{C}(\hat{\mathbf{S}})$, we perform the eigen-decomposition of $\hat{\mathbf{S}}\hat{\mathbf{S}}^{H}$ to obtain $\hat{\mathbf{U}}_{n}\in\mathbb{C}^{N\times(N-M)}$, which consists of the eigenvectors corresponding to the $N-M$ smallest eigenvalues. To proceed, we adopt the orthogonal projector $\hat{\boldsymbol{\Pi}}_{n}=\hat{\mathbf{U}}_{n}\hat{\mathbf{U}}_{n}^{H}$ [^5] We do not adopt $\hat{\boldsymbol{\Pi}}_{n}=\mathbf{I}_{N}-\hat{\mathbf{S}}(\hat{\mathbf{S}}^{H}\hat{\mathbf{S}})^{-1}\hat{\mathbf{S}}^{H}$ since numerical results show its low accuracy. onto the noise subspace estimate $\mathcal{C}(\hat{\mathbf{U}}_{n})$ and the Root-MUSIC algorithm [18] to obtain the DOA estimates $\hat{\theta}_{m}$’s.
Fig. 2 shows a scatter plot of $(\hat{\theta}_{1},\hat{\theta}_{2})$’s from both algorithms under 100100100 independent realizations. Here, $L=100$ and $K=K_{1}=1000$. The same snapshots of each realization are processed by both algorithms. We can know that under the same number of iterations $K=K_{1}$, both algorithms provide different $\hat{\mathbf{S}}$’s and for each realization, the $\hat{\mathbf{S}}\hat{\mathbf{S}}^{H}$ provided by the ECME algorithm is closer to the solution of $\mathbf{S}\mathbf{S}^{H}$ in (5). Hence, Fig. 3 compares the root mean square error (RMSE) performances of DOA estimation related to both algorithms with the Cramer–Rao lower bound (CRLB) [19, 20]. Here, $K=K_{1}=100$ and each RMSE is based on 100010001000 independent realizations. We observe that the RMSEs of DOA estimation from both algorithms approach the CRLBs as $L$ increases. More importantly, since the ECME algorithm yields faster convergence, its $\mathcal{C}(\mathcal{\hat{\mathbf{S}}})$ is more accurate under $K=K_{1}$ and hence its RMSE performance of DOA estimation is better.
Figure 2: Scatter plot of $(\hat{\theta}_{1},\hat{\theta}_{2})$’s from the FAAN method and the ECME algorithm.
Figure 3: RMSE performances of $(\hat{\theta}_{1},\hat{\theta}_{2})$ from the FAAN method and the ECME algorithm.
VI Conclusion
In this letter, the ECME algorithm, an extension of the EM algorithm, has been designed for MLFA, which has almost the same computational complexity at each iteration as the FAAN method with excellent convergence properties. However, numerical results have shown that the ECME algorithm yields faster stable convergence and is computationally more efficient. Importantly, the signal subspace estimated by the ECME algorithm can be used for the subspace based DOA estimation in unknown nonuniform noise.